# Attention Mechanisms Don’t Learn Additive Models: Rethinking Feature Importance for Transformers

### v1

@@ Appendix A Motivation: Failure Cases For Model Recovery
We create a dataset of 10 words (cf. Table 3) and train transformer models on samples from this dataset. We subsequently create sequences that repeatedly contain a single token $\tau$ (in this case, $\tau$=“perfect”), pass them through the transformers, and use Shapley values (approximated by Kernel-Shap) to explain their output. The result is visualized in Figure 6, and shows that a fully connected model (two-layer, 400 hidden units, ReLU) recovers the correct scores, whereas trainsformer models fail to reflect the true relationship. This shows that explanation methods that are explicitly or implicitly based on additive models lose their ability to recover the true data-generating process when transformer models are explained.
@@ B.1 Formalization of the transformer
Many popular LLMs follow the transformer architecture introduced by [55] with only minor modifications. We will introduce the most relevant building blocks of the architecture in this section. A schematic overview of the architecture is visualized in Figure 2. Let us denote the input embeddings for Layer $l\in 1,\ldots L$ by ${\bm{H}}^{(l-1)}=[{\bm{h}}^{(l-1)}_{1},\ldots,{\bm{h}}^{(l-1)}_{\lvert{\bm{t}}\rvert}]^{\top}\in\mathbb{R}^{{\lvert{\bm{t}}\rvert}\times d}$, where a single line ${\bm{h}}_{i}$ contains the embedding for token $i$. The input embeddings of the first layer consist of the token embeddings, i.e., ${\bm{H}}^{(0)}={\bm{E}}$. At the core of the architecture lies the attention head. For each token, a query, key, and a value vector are computed by applying an affine-linear transform. In matrix notation this can be written as
@@
where $\mathbf{1}_{{\lvert{\bm{t}}\rvert}}\in\mathbb{R}^{{\lvert{\bm{t}}\rvert}}$ denotes a vector of ones of length ${\lvert{\bm{t}}\rvert}$, ${\bm{b}}_{Q}^{(l)},{\bm{b}}_{V}^{(l)},{\bm{b}}_{K}^{(l)}\in\mathbb{R}^{d_{h}}$, ${\bm{W}}_{Q}^{(l)},{\bm{W}}_{K}^{(l)},{\bm{W}}_{V}^{(l)}\in\mathbb{R}^{d\times{d_{h}}},$ are trainable parameters and $d_{h}$ denotes the dimension of the attention head. [^1] We only formalize one attention head here, but consider the analogous case of multiple heads in our formal proofs. Keys and queries are projected onto each other and normalized by a row-wise softmax operation,
@@
This procedure is repeated iteratively for layers $1,\ldots,L$ such that we finally arrive at output embeddings ${\bm{H}}^{(L)}$. To perform classification, a classification head $\text{cls}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\lvert\mathcal{Y}\rvert}$ is put on top of a token at classification index $r$ (how this token is chosen depends on the architecture, common choices include $r\in\{1,{\lvert{\bm{t}}\rvert}\}$), such that we get the final logit output ${\bm{l}}=\text{cls}\left({\bm{h}}^{(L)}_{r}\right)$. The logit output is transformed to a probability vector via another softmax operation. Note that in the two-class case, we obtain the log-odds $F({\bm{t}})$ by taking the difference ($\Delta$) between logits
@@
Let $\mathcal{V}$ be a vocabulary and $C\geq 2,C\in\mathbb{N}$ be a maximum sequence length (context window). Let $w_{i}:\mathcal{V}\rightarrow\mathbb{R},\forall i\in 1,...,C$ be any mapping that assigns a token encountered at position $i$ a numerical score including at least one token $\tau\in\mathcal{V}$ with non-zero weight $w_{i}(\tau)\neq 0$ for some $i\in 2,\ldots,C$. Let $b\in\mathbb{R}$ be an arbitrary offset. Then, there exists no parametrization of the encoder or decoder single-layer transformer $F$ such that for every sequence ${\bm{t}}=[t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}]$ with length $1\leq{\lvert{\bm{t}}\rvert}\leq C$, the output of the transformer network is equivalent to
@@
as $\alpha_{rj}$ and $\bm{v}$ are independent of the token index $j$. We observe that the total input to final part $g$ is independent of $k$ in its entirety, as the first input ${\bm{e}}(\tau)$ is independent of $k$ and the second input is independent of $k$ as well. As $g$ is a deterministic function, also the log-odds output will be the same for all input sequences ${\bm{t}}_{k}$ and be independent of $k$. By the condition we have a non-zero weight $w_{j}(\tau)\neq 0$ for some $j\geq 2$. In this case, there are two sequences ${\bm{t}}_{j-1}$ (length $j{-}1$) and ${\bm{t}}_{j}$ (length $j$) consisting of only token $\tau$, where the outputs of the GAM follow
@@ B.3 Corollary: Transformers cannot represent linear models
Under the same conditions as in Proposition 4.1, a stack of multiple transformer blocks as in the model $F$ neither has a parametrization sufficient to represent the linear model.
@@
We show the result by induction, with the help of a lemma.
@@
For any mapping $s$, $v$ and a transformer with an embedding size $d,d_{h}\geq 3$, there exists a parameterization of the transformer to reflect the SLALOM model in Equation 4.
@@
We can prove the theorem by constructing a weight setup to reflect this mapping. We let the embedding ${\bm{e}}(\tau)$ be given by
@@
Multiple Layers. We can extend the argument to multiple layers by showing that the input vectors can just be forwarded by the transformer. This is simple by setting ${\bm{P}}\equiv\mathbf{0}$, the null-mapping, which can be represented by a linear operator. We then use the same classification hat as before. ∎
@@
Suppose query access to a model $G$ that takes sequences of tokens ${\bm{t}}$ with lengths $\lvert{\bm{t}}\rvert\in 1,\ldots,C$ and returns the log-odds according to a non-constant SLALOM on a vocabulary $\mathcal{V}$ with unknown parameter mappings $s:\mathcal{V}\rightarrow\mathbb{R}$, $v:\mathcal{V}\rightarrow\mathbb{R}$. For $C\geq 2$, we can recover the true mappings $s$, $v$ with $12\lvert\mathcal{V}\rvert-$ 1 queries (forward passes) of $F$.
@@
We can plug this back in to obtain the values for all token importance scores $s(\tau)=s(\theta)+\eta_{\theta}(\tau)$. We have thus computed the mappings $s$ and $v$ in $12\lvert\mathcal{V}\rvert-$ 1 forward passes, which completes the proof. ∎
@@
However, computing this sum remains usually intractable, as the number of coalitions grows exponentially. We can resort to common sampling approaches [10, 33] to approximate the sum.
Appendix C Additional Dicucssion and Intuition
@@ C.1 Generalization to multi-class problems
We can imagine the following generalizing SLALOM to multi class problems as follows: Suppose we have an importance mapping $s:\mathcal{V}\rightarrow\mathbb{R}$ that still maps each token to an importance score. However, we now introduce a value score mapping $v_{c}:\mathcal{V}\rightarrow\mathbb{R}$ for each class $c\in\mathcal{Y}$. Addtionally to requiring
@@ Appendix D Algorithm: Local SLALOM approximation
We propose the following algorithm to explain a sequence ${\bm{t}}=[t_{1},\ldots,t_{\lvert{\bm{t}}\rvert}]$ with SLALOM scores. In particular, we use the Mean-Squared-Error (MSE) to fit SLALOM on random sequences consisting of the individual tokens in the sequence. The procedure is given in Algorithm 1. To speed up the fitting we can sample a large collection of samples before the fitting loop and only sample minibatches from this collection in each step. We perform this optimization and use $b=5000$ samples in this work. We use sequences of length $n=2$.
@@
Algorithm 1 Local SLALOM approximation
@@ Appendix E Experimental Details
In this section, we provide details on the experimental setups. We provide the full source-code for the experimental evaluation in a ZIP file along with this submission to enhance reproducibility. We fully commit to open-sourcing our code in case of acceptance.
@@ E.1 Fitting transformers on a synthetic dataset
E.1.1 Dataset construction
We create a synthetic dataset to ensure a linear relationship between features and log-odds. Before sampling the dataset, we fix a vocabulary of tokens, ground truth scores for each token, and their occurrence probability. This means that each of the possible tokens already comes with a ground-truth score $w$ that has been manually assigned. The tokens, their respective scores $w$, and occurrence probabilities are listed in Table 3. Samples of the dataset are sampled in four steps that are exectued repeatedly for each sample:
@@ E.1.1 Dataset construction
    Having obtained the log-odds ratio for this sample F(t), we sample the labels according to this ratio. We have $p(y=1)/p(y=0)=\exp(F({\bm{t}}))$, which can be rearranged to $p(y=1)=\frac{\exp(F({\bm{t}}))}{1+\exp(F({\bm{t}}))}$. We sample a binary label $y$ for each sample according to this probability.
@@
E.1.2 Post-hoc fitting of surrogate models

### v2

@@ Appendix A Motivation: Failure Cases For Model Recovery
We create a dataset of 10 words (cf. Table 4) and train transformer models on samples from this dataset. We subsequently create sequences that repeatedly contain a single token $\tau$ (in this case, $\tau$=“perfect”), pass them through the transformers, and use Shapley values (approximated by Kernel-Shap) to explain their output. The result is visualized in Figure 7, and shows that a fully connected model (two-layer, 400 hidden units, ReLU) recovers the correct scores, whereas transformer models fail to reflect the true relationship. This shows that explanation methods that are explicitly or implicitly based on additive models lose their ability to recover the true data-generating process when transformer models are explained.
@@ B.1 Formalization of the transformer
Many popular LLMs follow the transformer architecture introduced by Vaswani et al. (2017) with only minor modifications. We will introduce the most relevant building blocks of the architecture in this section. A schematic overview of the architecture is visualized in Figure 2. Let us denote the input embeddings for Layer $l\in 1,\ldots L$ by ${\bm{H}}^{(l-1)}=[{\bm{h}}^{(l-1)}_{1},\ldots,{\bm{h}}^{(l-1)}_{\lvert{\bm{t}}\rvert}]^{\top}\in\mathbb{R}^{{\lvert{\bm{t}}\rvert}\times d}$, where a single line ${\bm{h}}_{i}$ contains the embedding for token $i$. The input embeddings of the first layer consist of the token embeddings, i.e., ${\bm{H}}^{(0)}={\bm{E}}$, where ${\bm{E}}=[{\bm{e}}_{1},\ldots,{\bm{e}}_{\lvert{\bm{t}}\rvert}]^{\top}\in\mathbb{R}^{{\lvert{\bm{t}}\rvert}\times d}$ is a matrix of the individual token embeddings. At the core of the architecture lies the attention head. For each token, a query, key, and a value vector are computed by applying an affine-linear transform. In matrix notation this can be written as
@@
where $\mathbf{1}_{{\lvert{\bm{t}}\rvert}}\in\mathbb{R}^{{\lvert{\bm{t}}\rvert}}$ denotes a vector of ones of length ${\lvert{\bm{t}}\rvert}$, ${\bm{b}}_{Q}^{(l)},{\bm{b}}_{V}^{(l)},{\bm{b}}_{K}^{(l)}\in\mathbb{R}^{d_{h}}$, ${\bm{W}}_{Q}^{(l)},{\bm{W}}_{K}^{(l)},{\bm{W}}_{V}^{(l)}\in\mathbb{R}^{d\times{d_{h}}},$ are trainable parameters and $d_{h}$ denotes the dimension of the attention head. [^2] We only formalize one attention head here, but consider the analogous caae of multiple heads in our formal proofs. Keys and queries are projected onto each other and normalized by a row-wise softmax operation,
@@
This procedure is repeated iteratively for layers $1,\ldots,L$ such that we finally arrive at output embeddings ${\bm{H}}^{(L)}$. To perform classification, a classification head $\text{cls}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\lvert\mathcal{Y}\rvert}$ is put on top of a token at classification index $r$ (how this token is chosen depends on the architecture, common choices include $r\in\{1,{\lvert{\bm{t}}\rvert}\}$), such that we get the final logit output ${\bm{l}}=\text{cls}\left({\bm{h}}^{(L)}_{r}\right)$. The logit output is transformed to a probability vector via another softmax operation. Note that in the two-class case, we obtain the log odds $F({\bm{t}})$ by taking the difference ($\Delta$) between logits
@@
Let $\mathcal{V}$ be a vocabulary and $C\geq 2,C\in\mathbb{N}$ be a maximum sequence length (context window). Let $w_{i}:\mathcal{V}\rightarrow\mathbb{R},\forall i\in 1,...,C$ be any map that assigns a token encountered at position $i$ a numerical score including at least one token $\tau\in\mathcal{V}$ with non-zero weight $w_{i}(\tau)\neq 0$ for some $i\in 2,\ldots,C$. Let $b\in\mathbb{R}$ be an arbitrary offset. Then, there exists no parametrization of the encoder or decoder single-layer transformer $F$ such that for every sequence ${\bm{t}}=[t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}]$ with length $1\leq{\lvert{\bm{t}}\rvert}\leq C$, the output of the transformer network is equivalent to
@@
as $\alpha_{rj}$ and $\bm{v}$ are independent of the token index $j$. We observe that the total input to final part $g$ is independent of $k$ in its entirety, as the first input ${\bm{e}}(\tau)$ is independent of $k$ and the second input is independent of $k$ as well. As $g$ is a deterministic function, also the log odds output will be the same for all input sequences ${\bm{t}}_{k}$ and be independent of $k$. By the condition we have a non-zero weight $w_{j}(\tau)\neq 0$ for some $j\geq 2$. In this case, there are two sequences ${\bm{t}}_{j-1}$ (length $j{-}1$) and ${\bm{t}}_{j}$ (length $j$) consisting of only token $\tau$, where the outputs of the additive model (GAM) follow
@@ B.3 Corollary: Transformers cannot represent linear models
Under the same conditions as in Proposition 4.1, a stack of multiple transformer blocks as in the model $F$ neither has a parametrization sufficient to represent the additive model.
@@
We show the result by induction with the help of a lemma.
@@
  • considering same-token sequences of variable length
@@
For any map $s$, $v$ and a transformer with an embedding size $d,d_{h}\geq 3$, there exists a parameterization of the transformer to reflect the SLALOM model in Equation 4.
@@
We can prove the theorem by constructing a weight setup to reflect this map. We let the embedding ${\bm{e}}(\tau)$ be given by
@@
Multiple Layers. We can extend the argument to multiple layers by showing that the input vectors can just be forwarded by the transformer. This is simple by setting ${\bm{P}}\equiv\mathbf{0}$, the null-map, which can be represented by a linear operator. We then use the same classification hat as before. ∎
@@
Suppose query access to a model $G$ that takes sequences of tokens ${\bm{t}}$ with lengths $\lvert{\bm{t}}\rvert\in 1,\ldots,C$ and returns the log odds according to a non-constant SLALOM on a vocabulary $\mathcal{V}$ with unknown parameter maps $s:\mathcal{V}\rightarrow\mathbb{R}$, $v:\mathcal{V}\rightarrow\mathbb{R}$. For $C\geq 2$, we can recover the true maps $s$, $v$ with $12\lvert\mathcal{V}\rvert-$ 1 queries (forward passes) of $F$.
@@
We can plug this back in to obtain the values for all token importance scores $s(\tau)=s(\theta)+\eta_{\theta}(\tau)$. We have thus computed the maps $s$ and $v$ in $12\lvert\mathcal{V}\rvert-$ 1 forward passes, which completes the proof. ∎
@@
However, computing this sum remains usually intractable, as the number of coalitions grows exponentially. We can resort to common sampling approaches (Castro et al., 2009; Maleki et al., 2013) to approximate the sum.
Appendix C Additional Discussion and Intuition
@@ C.1 Generalization to multi-class problems
We can imagine the following generalizing SLALOM to multi class problems as follows: We keep an importance map $s:\mathcal{V}\rightarrow\mathbb{R}$ that still maps each token to an importance score as previously. However, we now introduce a value score map $v_{c}:\mathcal{V}\rightarrow\mathbb{R}$ for each class $c\in\mathcal{Y}$. Additionally to requiring
@@
C.2 Practical Considerations
Our theoretical model contains slight deviations from real-world transformers to make it amendable to theoretical analysis. To represent token order, common architectures use positional embeddings, tying the embedding vectors to the token position $i$. The behavior that we show in this work’s analysis does however also govern transformers with positional embeddings for the following reason: While the positional embeddings could be used by the non-linear ffn part to differentiate sequences of different length in theory, our proofs show that to represent the linear model, the softmax operation must be inverted for any input sequence. This is a highly nonlinear operation and the number of possible sequences grows exponentially at a rate of $\lvert\mathcal{V}\rvert^{C}$ with the context length $C$. Learning-theoretic considerations (e.g., Bartlett et al., 1998) show that the number of input-output pairs the two-layer networks deployed can maximally represent is bounded by $\mathcal{O}\left(dn_{\text{hidden}}\log(dn_{\text{hidden}})\right)$, which is small ($d{=}786,n_{\text{hidden}}{=}3072$ for BERT) in contrast to the number of sequences ($C=1024$, $\lvert\mathcal{V}\rvert\approx 3\times 10^{4}$). We conclude that the inversion is therefore impossible for realistic setups and positional embeddings can be neglected, which is confirmed by our empirical findings.
Common models such as BERT also use a special token referred to as CLS-token where the classification head is placed on. In this work, we consider the CLS token just as a standard token in our analysis. In our empirical sections, we always append the CLS token as mandated by the architecture to make the sequences valid model inputs.
@@ Appendix D Algorithm: Local SLALOM approximation
We propose two algorithms to compute local explanations for a sequence ${\bm{t}}=[t_{1},\ldots,t_{\lvert{\bm{t}}\rvert}]$ with SLALOM scores. In particular, we use the Mean-Squared-Error (MSE) to fit SLALOM on modified sequences consisting of the individual tokens in the original sequence ${\bm{t}}$. To speed up the fitting we can sample a large collection of samples offline before the optimization.
D.1 Efficiently fitting SLALOM with SGD
For the efficient implementation SLALOM-eff given in Algorithm 1 we sample minibatches from this collection in each step and perform SGD steps on them. We perform this optimization using $b=5000$ samples in this work. We use sequences of $n=2$ random tokens from the sample for SLALOM-eff, making the forward passes through the model highly efficient.
D.2 Fitting SLALOM through iterative optimization
For the high-fidelity implementation SLALOM-fidel (Algorithm 2) we first use a different set of sequences and model scores to fit the surrogate: We delete up to $5$ tokens from the original sequence randomly to create the estimation dataset (similar to LIME). The fitting algorithm optimized for maximum fidelity uses an iterative optimization scheme to fit SLALOM models. It works by iteratively fitting ${\bm{v}}$ and ${\bm{s}}$ to the dataset obtained. Denote by ${\bm{f}}\in\mathbb{R}^{b}$ the model scores obtained for the $b$ input sequences ${\bm{t}}_{i},i=1...b$. As the SLALOM model in Equation 4 is a linear combination of the values score weighted by the normalized importance score, we can set up a matrix ${\bm{A}}$, where element ${\bm{a}}_{i,j}=\frac{\exp(s(t_{i}))}{\sum_{t_{j}\in{\bm{t}}_{i}}\exp(s(t_{j}))}$ provides the normalized importance for a given ${\bm{s}}$. We solve
$\min_{\bm{v}}({\bm{A}}{\bm{v}}-{\bm{f}})^{\top}({\bm{A}}{\bm{v}}-{\bm{f}}),$ (74)
for ${\bm{v}}$, which is a linear ordinary least squared problem that can be solved through the normal equation. This results in the optimal ${\bm{v}}$ for the given ${\bm{s}}$. In a second step, we keep ${\bm{v}}$ fixed and find better ${\bm{s}}$ scores. We can reformulate the equations for the samples as
$\sum_{t_{j}\in{\bm{t}}_{i}}\exp(s(t_{j}))v(t_{j})=\left(\sum_{t_{j}\in{\bm{t}}_{i}}\exp(s(t_{j}))\right){\bm{f}}_{i}$ (75)
$\Leftrightarrow\sum_{t_{j}\in{\bm{t}}_{i}}\underset{\bar{s}_{j}}{\underbrace{\exp(s(t_{j}))}}\underset{e_{i,j}}{\underbrace{(v(t_{j})-{\bm{f}}_{i})}}=0.$ (76)
This problem can be written with a vector $\bar{{\bm{s}}}\in\mathbb{R}^{|\mathcal{V}|}$ and a matrix ${\bm{E}}\in\mathbb{R}^{b\times|\mathcal{V}|}$ and results in an optimization problem
$\min_{\bar{{\bm{s}}}}$ $({\bm{E}}\bar{{\bm{s}}})^{\top}({\bm{E}}\bar{{\bm{s}}}),$ (77)
s.t. $\hat{{\bm{s}}}\geq\mathbf{0}$ (78)
$\lVert\hat{{\bm{s}}}\rVert_{1}\geq|\mathcal{V}|$ (79)
The conditions ensure that we can obtain the original ${\bm{s}}$-scores as $\log\hat{{\bm{s}}}$ (element-wise) and that the trivial solution $\hat{{\bm{s}}}=\mathbf{0}$ is not assumed. We solve this problem using a solver implemented in scipy.optimize.least_squares.
@@ D.2 Fitting SLALOM through iterative optimization
Algorithm 1 Local efficient SLALOM approximation (SLALOM-eff)
Sequence ${\bm{t}}$, trained model $F$ (outputs log odds), max. number of deletions $n$, learning rate $\lambda$, batch size $r$, sample pool size $b$, number of steps $c$
Initialize ${\bm{s}}=0$, ${\bm{s}}=0~{}~{}\forall~{}\text{unique}~{}t_{i}\in{\bm{t}}$
$B\leftarrow$ $b$ samples of random sequences of length $n$ obtained through deleting up to $n$ tokens randomly from ${\bm{t}}$.
$\text{steps}\leftarrow 0$
while $\text{steps}<c$ do
   ${\bm{v}}=\operatorname*{arg\,min}_{{\bm{v}}^{\prime}}\sum_{i=1}^{b}\left(F(B[i])-\text{SLALOM}_{{\bm{v}}^{\prime},{\bm{s}}}(B[i])\right)^{2}$ Fix ${\bm{s}}$ and optimize ${\bm{v}}$, OLS problem
   ${\bm{s}}=\operatorname*{arg\,min}_{{\bm{s}}^{\prime}}\sum_{i=1}^{b}\left(F(B[i])-\text{SLALOM}_{{\bm{v}},{\bm{s}}^{\prime}}(B[i])\right)^{2}$ Fix ${\bm{v}}$ and optimize ${\bm{s}}$, Quadratic problem
   $\text{steps}\leftarrow\text{steps}+1$
end while
return ${\bm{v}}$, ${\bm{s}}-\text{mean}({\bm{s}})$ normalize s to zero-mean
Algorithm 2 Local high-fidelity SLALOM approximation (SLALOM-fidel)
@@ Appendix E Experimental Details
In this section, we provide details on the experimental setups. We provide the full source-code in our GitHub repository.
@@ E.1 Fitting transformers on a synthetic dataset
E.1.1 Dataset construction: Linear Dataset
We create a synthetic dataset to ensure a linear relationship between features and log odds. Before sampling the dataset, we fix a vocabulary of tokens, ground truth scores for each token, and their occurrence probability. This means that each of the possible tokens already comes with a ground-truth score $w$ that has been manually assigned. The tokens, their respective scores $w$, and occurrence probabilities are listed in Table 4. Samples of the dataset are sampled in four steps that are exectued repeatedly for each sample:
@@ E.1.1 Dataset construction: Linear Dataset
    Having obtained the log odds ratio for this sample F(t), we sample the labels according to this ratio. We have $p(y=1)/p(y=0)=\exp(F({\bm{t}}))$, which can be rearranged to $p(y=1)=\frac{\exp(F({\bm{t}}))}{1+\exp(F({\bm{t}}))}$. We sample a binary label $y$ for each sample according to this probability.
@@
E.1.2 Dataset Construction: SLALOM dataset
We resort to a second synthetic dataset to study the recovery property for the SLALOM. This required a SLALOM relation between the data features and the labels. To find a realistic distribution of scores, we compute a BoW importance scores for input tokens of the BERT model on the IMDB dataset by counting the class-wise occurrence probabilities. We select $200$ tokens randomly from this dataset. We use these scores as value scores $v$ but multiply them by a factors of $2$ as many words have very small BoW importances. In realistic datasets, we observed that value scores $v$ are correlated with the importance scores $s$. Therefore, we sample
which results in the value/importance distribution given in Figure 8. We assign each word an equal occurrence probability and sample sequences of words at uniformly distributed lengths in random $[1,30]$. After a sequence is sampled, labels are subsequently sampled according to the log odds ratio of the SLALOM.
E.2 Post-hoc fitting of surrogate models